A random variable X has the probability density function given below

g(x)= {2x,0≤x<1
{ 6-x,1≤x<2
{ 0, otherwise

Show graphically and find out the probabilities: P(X≤3)
1
Expert's answer
2020-08-25T11:03:29-0400

"g(x) = \begin{cases} 2x, & 0 \leq x<1 \\ 6-x, & 1\leq x <2 \\ 0, & otherwise\end{cases}"

Plotting this, we get following



However, in order to be indeed probability density function condition "\int_{-\infty}^{+\infty} g(x)dx =1" must be fulfilled. In our case,

"\int_0^1(2x)dx + \int_1^2(6-x)dx = x^2|_0^1 + 6x|_1^2 - \frac{x^2}{2}|_1^2 = 1 +6 -2 - 0.5 = 4.5."

In such case, g(x) should be normalized, namely "\tilde{g}(x) = g(x)/4.5".

In order to find "P(X \leq3)" we need to integrate "\tilde{g}(x)" from "-\infty" to 3 or, what is the same, find the area below the (normalized) plot. Since for "x \geq 3" "\tilde{g}(x) =0", integration from "-\infty" to 3 is the same as integration from "-\infty" to "+\infty". The last one we already did above.

"P(X \leq 3) = \int_{-\infty}^3 g(x)dx =\frac{\int_{-\infty}^3 \tilde{g}(x)dx }{4.5} = \frac{\int_{-\infty}^{+\infty}\tilde{g}(x)dx }{4.5} = 1".

Alternative approach: "P(X \leq 3) = 1 - P(X \geq 3) = 1", since for "x \geq 3" "g(x) =0" and integral of it is 0 too.

Answer: "P(X \leq 3) = 1."